\begin{thebibliography}{10}

\bibitem{PRD.1997.55.6760}
D.~Colladay and V.~A. Kosteleck\'y.
\newblock $\mathrm{CPT}$ violation and the standard model.
\newblock {\em Phys. Rev. D}, 55:6760, Jun 1997.

\bibitem{PRD.1998.58.116002}
D.~Colladay and V.~A. Kosteleck\'y.
\newblock Lorentz-violating extension of the standard model.
\newblock {\em Phys. Rev. D}, 58:116002, Oct 1998.

\bibitem{PRD.2004.69.105009}
V.~A. Kosteleck\'y.
\newblock Gravity, lorentz violation, and the standard model.
\newblock {\em Phys. Rev. D}, 69:105009, May 2004.

\bibitem{PRD.1999.59.116008}
S.~Coleman and S.~L. Glashow.
\newblock High-energy tests of lorentz invariance.
\newblock {\em Phys. Rev. D}, 59:116008, Apr 1999.

\bibitem{PRL.1989.63.224}
V.~A. Kosteleck\'y and S.~Samuel.
\newblock Phenomenological gravitational constraints on strings and
  higher-dimensional theories.
\newblock {\em Phys. Rev. Lett.}, 63:224, Jul 1989.

\bibitem{PRL.1991.66.1811}
V.~A. Kosteleck\'y and S.~Samuel.
\newblock Photon and graviton masses in string theories.
\newblock {\em Phys. Rev. Lett.}, 66:1811, Apr 1991.

\bibitem{PRL.1999.82.3572}
R.~Jackiw and V.~A. Kosteleck\'{y}.
\newblock Radiatively induced lorentz and $\mathit{CPT}$ violation in
  electrodynamics.
\newblock {\em Phys. Rev. Lett.}, 82:3572, May 1999.

\bibitem{PRL.1999.83.2518}
M.~P\'{e}rez-Victoria.
\newblock Exact calculation of the radiatively induced lorentz and
  $\mathit{CPT}$ violation in qed.
\newblock {\em Phys. Rev. Lett.}, 83:2518, Sep 1999.

\bibitem{PRD.1999.60.127901}
J.~M. Chung.
\newblock Lorentz- and \textit{CPT} -violating chern-simons term in the
  functional integral formalism.
\newblock {\em Phys. Rev. D}, 60:127901, Oct 1999.

\bibitem{PRD.2001.63.105015}
J.~M. Chung and B.~K. Chung.
\newblock Induced lorentz- and $\mathrm{CPT}$-violating chern-simons term in
  qed: Fock-schwinger proper time method.
\newblock {\em Phys. Rev. D}, 63:105015, Apr 2001.

\bibitem{PRD.2001.64.046013}
A.~P.~B. Scarpelli, M.~Sampaio, M.~C. Nemes, and B.~Hiller.
\newblock Chiral anomaly and $\mathrm{CPT}$ invariance in an implicit momentum
  space regularization framework.
\newblock {\em Phys. Rev. D}, 64:046013, Jul 2001.

\bibitem{JPA.2003.36.4937}
D.~Bazeia, T.~Mariz, J.~R. Nascimento, E.~Passos, and R.~F. Ribeiro.
\newblock Lorentz and cpt symmetries in commutative and noncommutative
  spacetimes.
\newblock {\em J. Phys. A}, 36(17):4937, 2003.

\bibitem{PRD.2006.73.65015}
M.~N. Barreto, D.~Bazeia, and R.~Menezes.
\newblock Defect structures in lorentz and cpt violating scenarios.
\newblock {\em Phys. Rev. D}, 73(6):065015--, March 2006.

\bibitem{PRD.2009.79.123503}
P.~P. Avelino, D.~Bazeia, L.~Losano, R.~Menezes, and J.~J. Rodrigues.
\newblock Impact of lorentz violation on the dynamics of inflation.
\newblock {\em Phys. Rev. D}, 79(12):123503, June 2009.

\bibitem{PD.2010.239.942}
D.~Bazeia, M.~M Ferreira~Jr, A~R Gomes, and R~Menezes.
\newblock Lorentz-violating effects on topological defects generated by two
  real scalar fields.
\newblock {\em Phys. D}, 239(12):942, 2010.

\bibitem{PRD.2012.86.065011}
C.~Miller, R.~Casana, M.~M. Ferreira~Jr., and E.~da~Hora.
\newblock Uncharged compactlike and fractional lorentz-violating bps vortices
  in the $cpt$-even sector of the standard model extension.
\newblock {\em Phys. Rev. D}, 86:065011, Sep 2012.

\bibitem{PRD.2008.78.125013}
R.~Casana, M.~M. Ferreira~Jr., and J.~S. Rodrigues.
\newblock Lorentz-violating contributions of the carroll-field-jackiw model to
  the cmb anisotropy.
\newblock {\em Phys. Rev. D}, 78:125013, Dec 2008.

\bibitem{PRD.2011.84.076006}
B.~Altschul.
\newblock Bounding lorentz violation at particle colliders by tracking the
  motion of charged particles.
\newblock {\em Phys. Rev. D}, 84:076006, Oct 2011.

\bibitem{EPL.2011.96.61001}
O.~Ganguly, D.~Gangopadhyay, and P~Majumdar.
\newblock Lorentz-preserving fields in lorentz-violating theories.
\newblock {\em Eur. Phys. Lett.}, 96(6):61001, 2011.

\bibitem{PRD.2012.85.085023}
M.~Cambiaso, R.~Lehnert, and R.~Potting.
\newblock Massive photons and lorentz violation.
\newblock {\em Phys. Rev. D}, 85:085023, Apr 2012.

\bibitem{PRD.2012.85.105001}
M.~Pospelov and Y.~Shang.
\newblock Lorentz violation in ho\ifmmode \check{r}\else
  \v{r}\fi{}ava-lifshitz-type theories.
\newblock {\em Phys. Rev. D}, 85:105001, May 2012.

\bibitem{EPL.2012.99.21003}
J.~Leite and T.~Mariz.
\newblock Induced lorentz-violating terms at finite temperature.
\newblock {\em Eur. Phys. Lett.}, 99(2):21003, 2012.

\bibitem{PRD.2011.84.045008}
R.~Casana, E.~S. Carvalho, and M.~M. Ferreira~Jr.
\newblock Dimensional reduction of the $cpt$-even electromagnetic sector of the
  standard model extension.
\newblock {\em Phys. Rev. D}, 84:045008, Aug 2011.

\bibitem{arXiv:1304.2090}
R.~V. Maluf, V.~Santos, W.~T. Cruz, and C.~A.~S. Almeida.
\newblock Matter-gravity scattering in the presence of spontaneous lorentz
  violation.
\newblock {\em arXiv:1304.2090 [hep-th]}, 2013.

\bibitem{NPB.2003.657.214}
C.~Adam and F.~R. Klinkhamer.
\newblock Photon decay in a cpt-violating extension of quantum electrodynamics.
\newblock {\em Nucl. Phys. B}, 657(0):214, May 2003.

\bibitem{NPB.2001.607.247}
C.~Adam and F.~R. Klinkhamer.
\newblock Causality and cpt violation from an abelian chern-simons-like term.
\newblock {\em Nucl. Phys. B}, 607(1-2):247, July 2001.

\bibitem{PRD.1995.51.5961}
A.~A. Andrianov and R.~Soldati.
\newblock Lorentz symmetry breaking in abelian vector-field models with
  wess-zumino interaction.
\newblock {\em Phys. Rev. D}, 51(10):5961, May 1995.

\bibitem{PRD.1998.59.25002}
A.~A. Andrianov, R.~Soldati, and L.~Sorbo.
\newblock Dynamical lorentz symmetry breaking from a (3+1)-dimensional
  axion-wess-zumino model.
\newblock {\em Phys. Rev. D}, 59(2):025002, December 1998.

\bibitem{PLB.1998.435.449}
A.~A. Andrianov and R.~Soldati.
\newblock Patterns of lorentz symmetry breaking in qed by cpt-odd interaction.
\newblock {\em Phys. Lett. B}, 435(3-4):449, September 1998.

\bibitem{PRD.2003.67.125011}
H.~Belich, M.~M. Ferreira~Jr., J.~A. Helay\"el-Neto, and M.~T.~D. Orlando.
\newblock Dimensional reduction of a lorentz- and \textit{CPT} -violating
  maxwell-chern-simons model.
\newblock {\em Phys. Rev. D}, 67:125011, Jun 2003.

\bibitem{PRD.2009.80.125040}
R.~Casana, M.~M. Ferreira~Jr., A.~R. Gomes, and P.~R.~D. Pinheiro.
\newblock Gauge propagator and physical consistency of the $cpt$-even part of
  the standard model extension.
\newblock {\em Phys. Rev. D}, 80:125040, Dec 2009.

\bibitem{EPJC.2008.56.571}
A.~P.~Baêta Scarpelli, Marcos Sampaio, M.~C. Nemes, and B.~Hiller.
\newblock Gauge invariance and the cpt and lorentz violating induced
  chern–simons-like term in extended qed.
\newblock {\em Eur. Phys. J. C}, 56(4):571, 2008.

\bibitem{PRD.2010.81.105015}
R.~Casana, M.~M. Ferreira~Jr., and M.~R.~O. Silva.
\newblock Parity-odd and $cpt$-even electrodynamics of the standard model
  extension at finite temperature.
\newblock {\em Phys. Rev. D}, 81:105015, May 2010.

\bibitem{PRD.2011.83.045018}
T.~Mariz.
\newblock Radiatively induced lorentz-violating operator of mass dimension five
  in qed.
\newblock {\em Phys. Rev. D}, 83:045018, Feb 2011.

\bibitem{EPJC.2012.72.2070}
R.~Casana, M.~M. Ferreira~Jr., and R.~P.~M. Moreira.
\newblock Consistency analysis of a nonbirefringent lorentz-violating planar
  model.
\newblock {\em Eur. Phys. J. C}, 72:2070, 2012.

\bibitem{PRD.2012.86.045003}
B.~Charneski, M.~Gomes, R.~V. Maluf, and A.~J. da~Silva.
\newblock Lorentz violation bounds on bhabha scattering.
\newblock {\em Phys. Rev. D}, 86:045003, Aug 2012.

\bibitem{JPG.2012.39.125001}
A.~P.~B. Scarpelli.
\newblock Qed with chiral nonminimal coupling: aspects of the lorentz-violating
  quantum corrections.
\newblock {\em J. Phys. G}, 39(12):125001, 2012.

\bibitem{JPG.2012.39.35002}
G.~Gazzola, H.~G. Fargnoli, A.~P.~Ba\^{e}ta Scarpelli, M.~Sampaio, and M.~C.
  Nemes.
\newblock Qed with minimal and nonminimal couplings: on the quantum generation
  of lorentz-violating terms in the pure photon sector.
\newblock {\em J. Phys. G}, 39(3):035002, 2012.

\bibitem{PRD.2002.66.081302}
T.~Jacobson, S.~Liberati, and D.~Mattingly.
\newblock Tev astrophysics constraints on planck scale lorentz violation.
\newblock {\em Phys. Rev. D}, 66:081302, Oct 2002.

\bibitem{AR.2009.59.245}
S.~Liberati and L.~Maccione.
\newblock Lorentz violation: Motivation and new constraints.
\newblock {\em Ann. Rev. Nuclear Particle Science}, 59:245, 2009.

\bibitem{PRD.2011.83.127702}
L.~Shao and B.~Ma.
\newblock Lorentz-violation-induced vacuum birefringence and its astrophysical
  consequences.
\newblock {\em Phys. Rev. D}, 83:127702, Jun 2011.

\bibitem{RMP.2011.83.11}
V.~A. Kosteleck\'{y} and N.~Russell.
\newblock Data tables for lorentz and cpt violation.
\newblock {\em Rev. Mod. Phys.}, 83(1):11, March 2011.

\bibitem{PRL.2001.87.251304}
V.~A. Kosteleck\'y and M.~Mewes.
\newblock Cosmological constraints on lorentz violation in electrodynamics.
\newblock {\em Phys. Rev. Lett.}, 87:251304, Nov 2001.

\bibitem{PRD.2002.66.056005}
V.~A. Kosteleck\'y and M.~Mewes.
\newblock Signals for lorentz violation in electrodynamics.
\newblock {\em Phys. Rev. D}, 66:056005, Sep 2002.

\bibitem{PRL.1997.79.1432}
R.~Bluhm, V.~A. Kosteleck\'y, and N.~Russell.
\newblock Testing $\mathit{CPT}$ with anomalous magnetic moments.
\newblock {\em Phys. Rev. Lett.}, 79:1432, Aug 1997.

\bibitem{PRD.1998.57.3932}
R.~Bluhm, V.~A. Kosteleck\'y, and N.~Russell.
\newblock $\mathrm{CPT}$ and lorentz tests in penning traps.
\newblock {\em Phys. Rev. D}, 57:3932, Apr 1998.

\bibitem{PRL.2002.88.090801}
R.~Bluhm, V.~A. Kosteleck\'y, C.~D. Lane, and N.~Russell.
\newblock Clock-comparison tests of lorentz and $\mathit{CPT}$ symmetry in
  space.
\newblock {\em Phys. Rev. Lett.}, 88:090801, Feb 2002.

\bibitem{PRL.2000.84.1381}
R.~Bluhm and V.~A. Kosteleck\'y.
\newblock Lorentz and $\mathit{CPT}$ tests with spin-polarized solids.
\newblock {\em Phys. Rev. Lett.}, 84:1381, Feb 2000.

\bibitem{PRL.2000.84.1098}
R.~Bluhm, V.~A. Kosteleck\'y, and C.~D. Lane.
\newblock $\mathit{CPT}$ and lorentz tests with muons.
\newblock {\em Phys. Rev. Lett.}, 84:1098, Feb 2000.

\bibitem{PRL.1999.82.2254}
R.~Bluhm, V.~A. Kosteleck\'y, and N.~Russell.
\newblock $\mathit{CPT}$ and lorentz tests in hydrogen and antihydrogen.
\newblock {\em Phys. Rev. Lett.}, 82:2254, Mar 1999.

\bibitem{EPJC.2005.41.421}
H.~Belich, T.~Costa-Soares, M.~M. Ferreira~Jr, and J.~A. Helayël-Neto.
\newblock Non-minimal coupling to a lorentz-violating background and
  topological implications.
\newblock {\em Eur. Phys. J. C}, 41:421, 2005.

\bibitem{PRL.1984.53.319}
Y.~Aharonov and A.~Casher.
\newblock Topological quantum effects for neutral particles.
\newblock {\em Phys. Rev. Lett.}, 53(4):319, July 1984.

\bibitem{PRD.2012.86.045001}
A.~G. Grushin.
\newblock Consequences of a condensed matter realization of lorentz-violating
  qed in weyl semi-metals.
\newblock {\em Phys. Rev. D}, 86:045001, Aug 2012.

\bibitem{ADP.2011.523.910}
K.~Bakke, H.~Belich, and E.~O. Silva.
\newblock Relativistic anandan quantum phase in the lorentz violation
  background.
\newblock {\em Ann. Phys. (Berlin)}, 523(11):910, 2011.

\bibitem{JMP.2011.52.063505}
K.~Bakke, H.~Belich, and E.~O. Silva.
\newblock Relativistic landau-aharonov-casher quantization based on the lorentz
  symmetry violation background.
\newblock {\em J. Math. Phys.}, 52:063505, 2011.

\bibitem{PRD.2011.83.125025}
H.~Belich, E.~O. Silva, M.~M. Ferreira~Jr, and M.~T.~D. Orlando.
\newblock Aharonov-bohm-casher problem with a nonminimal lorentz-violating
  coupling.
\newblock {\em Phys. Rev. D}, 83(12):125025, June 2011.

\bibitem{PLB.2006.639.675}
H.~Belich, T.~Costa-Soares, M.~M. Ferreira~Jr, J.~A. Helay\"{e}l-Neto, and
  M.~T.~D Orlando.
\newblock A comment on the topological phase for anti-particles in a
  lorentz-violating environment.
\newblock {\em Phys. Lett. B}, 639(6):675, 2006.

\bibitem{EPJC.2009.62.425}
H.~Belich, L.~P. Colatto, T.~Costa-Soares, J.~A. Helay\"{e}l-Neto, and M.~T.~D.
  Orlando.
\newblock Magnetic moment generation from non-minimal couplings in a scenario
  with lorentz-symmetry violation.
\newblock {\em Eur. Phys. J. C}, 62:425, 2009.

\bibitem{JPG.2012.39.055004}
K.~Bakke, E.~O. Silva, and H.~Belich.
\newblock He–mckellar–wilkens effect and scalar aharonov–bohm effect for
  a neutral particle based on the lorentz symmetry violation.
\newblock {\em J. Phys. G}, 39(5):055004, 2012.

\bibitem{JPG.2012.39.105004}
L.~R. Ribeiro, C.~Furtado, and E.~Passos.
\newblock An analogy of the quantum hall conductivity in a lorentz-symmetry
  violation setup.
\newblock {\em J. Phys. G}, 39(10):105004, 2012.

\bibitem{AP.2013.333.272}
K.~Bakke and H.~Belich.
\newblock Relativistic landau-he-mckellar-wilkens quantization and relativistic
  bound states solutions for a coulomb-like potential induced by the lorentz
  symmetry breaking effects.
\newblock {\em Ann. Phys. (N.Y.)}, 333(0):272, 2013.

\bibitem{JPG.2012.39.085001}
K~Bakke and H~Belich.
\newblock Abelian geometric phase for a dirac neutral particle in a lorentz
  symmetry violation environment.
\newblock {\em J. Phys. G}, 39(8):085001, 2012.

\bibitem{JPG.2013.40.065002}
K~Bakke and H~Belich.
\newblock Quantum holonomies based on the lorentz-violating tensor background.
\newblock {\em J. Phys. G}, 40(6):065002, 2013.

\bibitem{PRD.2013.87.047701}
R.~Casana, M.~M. Ferreira~Jr., E.~Passos, F.~E.~P. dos Santos, and E.~O. Silva.
\newblock New cpt-even and lorentz-violating nonminimal coupling in the dirac
  equation.
\newblock {\em Phys. Rev. D}, 87(4):047701--, February 2013.

\bibitem{PRD.2012.86.125033}
R.~Casana, M.~M. Ferreira~Jr., R.~V. Maluf, and F.~E.~P. dos Santos.
\newblock Effects of a $cpt$-even and lorentz-violating nonminimal coupling on
  electron-positron scattering.
\newblock {\em Phys. Rev. D}, 86:125033, Dec 2012.

\bibitem{PRD.1978.18.2932}
H.~J. de~Vega.
\newblock Fermions and vortex solutions in abelian and non-abelian gauge
  theories.
\newblock {\em Phys. Rev. D}, 18(8):2932--2944, October 1978.

\bibitem{NPB.1988.307.909}
R.~H. Brandenberger, A.~Davis, and A.~M. Matheson.
\newblock Callan-rubakov effect for strings.
\newblock {\em Nucl. Phys. B}, 307(4):909--923, October 1988.

\bibitem{PRL.1989.62.1071}
M.~G. Alford and F.~Wilczek.
\newblock Aharonov-bohm interaction of cosmic strings with matter.
\newblock {\em Phys. Rev. Lett.}, 62(10):1071--1074, March 1989.

\bibitem{EPJC.2013.73.2402}
E.~O. Silva, F.~M. Andrade, C.~Filgueiras, and H.~Belich.
\newblock On aharonov-casher bound states.
\newblock {\em Eur. Phys. J. C}, 73(4):2402, 2013.

\bibitem{JMP.1985.26.2520}
W.~Bulla and F.~Gesztesy.
\newblock Deficiency indices and singular boundary conditions in quantum
  mechanics.
\newblock {\em J. Math. Phys.}, 26(10):2520, 1985.

\bibitem{Book.2004.Albeverio}
S.~Albeverio, F.~Gesztesy, R.~Hoegh-Krohn, and H.~Holden.
\newblock {\em Solvable Models in Quantum Mechanics}.
\newblock AMS Chelsea Publishing, Providence, RI, second edition, 2004.

\bibitem{Book.1975.Reed.II}
M.~Reed and B.~Simon.
\newblock {\em Methods of Modern Mathematical Physics. II. Fourier Analysis,
  Self-Adjointness.}
\newblock Academic Press, New York - London, 1975.

\bibitem{Book.2011.Sakurai}
J.~J. Sakurai and J.~Napolitano.
\newblock {\em Modern Quantum Mechanics}.
\newblock Addison-Wesley, 2nd ed. edition, 2011.

\bibitem{JPA.1995.28.2359}
J.~Audretsch, U.~Jasper, and V.~D. Skarzhinsky.
\newblock A pragmatic approach to the problem of the self-adjoint extension of
  hamilton operators with the aharonov-bohm potential.
\newblock {\em J. Phys. A}, 28(8):2359, 1995.

\bibitem{PRA.1992.46.6052}
F.~A.~B. Coutinho, Y.~Nogami, and J.~Fernando~Perez.
\newblock Self-adjoint extensions of the hamiltonian for a charged particle in
  the presence of a thread of magnetic flux.
\newblock {\em Phys. Rev. A}, 46(9):6052, November 1992.

\bibitem{JPA.2010.43.354011}
C.~R. de~Oliveira and M.~Pereira.
\newblock Scattering and self-adjoint extensions of the aharonov–bohm
  hamiltonian.
\newblock {\em J. Phys. A}, 43(35):354011, 2010.

\bibitem{Book.1972.Abramowitz}
M.~Abramowitz and I.~A. Stegun, editors.
\newblock {\em Handbook of Mathematical Functions}.
\newblock New York: Dover Publications, 1972.

\bibitem{AP.1983.146.1}
S.~N.~M. Ruijsenaars.
\newblock The aharonov-bohm effect and scattering theory.
\newblock {\em Ann. Phys. (NY)}, 146(1):1, March 1983.

\bibitem{PRC.1999.60.34308}
K.~Bennaceur, J.~Dobaczewski, and M.~Ploszajczak.
\newblock Continuum effects for the mean-field and pairing properties of weakly
  bound nuclei.
\newblock {\em Phys. Rev. C}, 60(3):034308, August 1999.

\bibitem{EPL.2013.101.51005}
E.~O. Silva and F.~M. Andrade.
\newblock Remarks on the aharonov-casher dynamics in a cpt-odd
  lorentz-violating background.
\newblock {\em Europhys. Lett.}, 101(5):51005--, 2013.

\bibitem{AP.2010.325.2529}
C.~Filgueiras, E.~O. Silva, W.~Oliveira, and F.~Moraes.
\newblock The effect of singular potentials on the harmonic oscillator.
\newblock {\em Ann. Phys. (NY)}, 325(11):2529, November 2010.

\bibitem{PLB.2013.719.467}
F.~M. Andrade and E.~O. Silva.
\newblock Effects of quantum deformation on the spin-1/2 aharonov-bohm problem.
\newblock {\em Phys. Lett. B}, 719(4-5):467, February 2013.

\bibitem{JMP.2012.53.122106}
C.~Filgueiras, E.~O. Silva, and F.~M. Andrade.
\newblock Nonrelativistic quantum dynamics on a cone with and without a
  constraining potential.
\newblock {\em J. Math. Phys.}, 53(12):122106, December 2012.

\bibitem{PLB.1994.333.238}
M.~Bordag and S.~Voropaev.
\newblock Bound states of an electron in the field of the magnetic string.
\newblock {\em Phys. Lett. B}, 333(1–2):238, July 1994.

\bibitem{PRL.1990.64.2347}
C.~R. Hagen.
\newblock Exact equivalence of spin-1/2 aharonov-bohm and aharonov-casher
  effects.
\newblock {\em Phys. Rev. Lett.}, 64(20):2347, May 1990.

\bibitem{PRD.1977.16.1815}
A.~S. Goldhaber.
\newblock Dirac particle in a magnetic field: Symmetries and their breaking by
  monopole singularities.
\newblock {\em Phys. Rev. D}, 16(6):1815, September 1977.

\bibitem{PRD.1977.15.2287}
Y.~Kazama, C.~N. Yang, and A.~S. Goldhaber.
\newblock Scattering of a dirac particle with charge z e by a fixed magnetic
  monopole.
\newblock {\em Phys. Rev. D}, 15(8):2287, April 1977.

\bibitem{PRL.1983.50.464}
B.~Grossman.
\newblock Does a dyon leak?
\newblock {\em Phys. Rev. Lett.}, 50(7):464, February 1983.

\bibitem{PLB.1993.298.63}
N.~Ganoulis.
\newblock Modified aharonov-bohm scattering from electroweak strings.
\newblock {\em Phys. Lett. B}, 298(1–2):63, January 1993.

\bibitem{NPB.1994.419.323}
A.~C. Davis, A.~P. Martin, and N.~Ganoulis.
\newblock Charged-particle scattering from electroweak and semi-local strings.
\newblock {\em Nucl. Phys. B}, 419(2):323, May 1994.

\bibitem{PRD.1994.49.2092}
F.~A.~B. Coutinho and J.~F. Perez.
\newblock Helicity conservation in the aharonov-bohm scattering of dirac
  particles.
\newblock {\em Phys. Rev. D}, 49(4):2092, February 1994.

\bibitem{JPA.2001.34.8859}
V.~S. Araujo, F.~A.~B. Coutinho, and J.~F. Perez.
\newblock On the most general boundary conditions for the aharonov-bohm
  scattering of a dirac particle: helicity and aharonov-bohm symmetry
  conservation.
\newblock {\em J. Phys. A}, 34(42):8859, 2001.

\bibitem{AJP.2001.69.322}
G.~Bonneau, J.~Faraut, and G.~Valent.
\newblock Self-adjoint extensions of operators and the teaching of quantum
  mechanics.
\newblock {\em Am. J. Phys.}, 69(3):322, March 2001.

\end{thebibliography}

\documentclass[12pt,a4paper]{iopart}

\expandafter\let\csname equation*\endcsname\relax
\expandafter\let\csname endequation*\endcsname\relax
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{hyperref}
\usepackage{iopams}

\RequirePackage{color}

\hypersetup{
  colorlinks=true,linkcolor=blue,citecolor=blue,
  filecolor=blue,urlcolor=blue,breaklinks=true
}

\begin{document}

\title{On the Aharonov-Casher scattering in a CPT-odd
  Lorentz-violating background}

\author{F. M. Andrade}
\ead{fmandrade@uepg.br}
\address{
  Departamento de Matem\'{a}tica e Estat\'{i}stica,
  Universidade Estadual de Ponta Grossa,
  84030-900 Ponta Grossa-PR, Brazil
}

\author{E. O. Silva}
\ead{edilbertoo@gmail.com}
\address{
  Departamento de F\'{i}sica,
  Universidade Federal do Maranh\~{a}o,
  Campus Universit\'{a}rio do Bacanga,
  65085-580 S\~{a}o Lu\'{i}s-MA, Brazil
}
\author{T. Prud\^{e}ncio}
\ead{thprudencio@gmail.com}
\address{
  Instituto de F\'{i}sica,
  Universidade de Bras\'{i}lia,
  Caixa Postal 04455,
  70919-970, Bras\'ilia-DF, Brazil.
}

\author{C. Filgueiras}
\ead{cleversonfilgueiras@yahoo.com.br}
\address{
  Departamento de F\'{i}sica,
  Universidade Federal de Campina Grande,
  Caixa Postal 10071,
  58109-970 Campina Grande-PB, Brazil
}

\date{\today}

\begin{abstract}
The effects of a Lorentz symmetry violating background vector on
the Aharonov-Casher scattering in the nonrelativistic limit
is considered.
By using the self-adjoint extension method we found that there
is an additional scattering for any value of the self-adjoint
extension parameter and non-zero energy bound states for
negative values of this parameter.
Expressions for the energy bound states, phase-shift and the
scattering matrix are explicitly determined in terms of the
self-adjoint extension parameter.
The expression obtained for the scattering amplitude reveals
that the helicity is not conserved in this scenario.
\end{abstract}

\pacs{11.30.Cp, 03.65.Ge, 03.65.Db, 11.55.-m}
\submitto{\jpg}
\maketitle

\section{Introduction}
\label{sec:introduction}

The standard model extension (SME) proposed by Colladay and
Kosteleck\'{y}
\cite{PRD.1997.55.6760,PRD.1998.58.116002,PRD.2004.69.105009}
(cf. also Refs.
\cite{PRD.1999.59.116008,PRL.1989.63.224,PRL.1991.66.1811}) has
been an usual framework for investigating signals of Lorentz
violation in physical systems and has inspired a great deal of
investigations in this theme in recent years.
The interest in this issue appears in the different contexts,
such as field theory
 \cite{PRL.1999.82.3572,PRL.1999.83.2518,PRD.1999.60.127901,
PRD.2001.63.105015,PRD.2001.64.046013,JPA.2003.36.4937,
PRD.2006.73.65015,PRD.2009.79.123503,PD.2010.239.942,
PRD.2012.86.065011,PRD.2008.78.125013,PRD.2011.84.076006,
EPL.2011.96.61001,PRD.2012.85.085023,PRD.2012.85.105001,
EPL.2012.99.21003,PRD.2011.84.045008,arXiv:1304.2090}
aspects on the gauge sector of the SME
\cite{NPB.2003.657.214,NPB.2001.607.247,PRD.1995.51.5961,
PRD.1998.59.25002,PLB.1998.435.449,PRD.2003.67.125011,
PRD.2009.80.125040},
quantum electrodynamics
\cite{EPJC.2008.56.571,PRD.2010.81.105015,PRD.2011.83.045018,
EPJC.2012.72.2070,PRD.2012.86.045003,JPG.2012.39.125001,JPG.2012.39.35002},
and astrophysics
\cite{PRD.2002.66.081302,AR.2009.59.245,PRD.2011.83.127702}.
These many contributions have elucidated effects induced by
Lorentz violation and served to set up stringent upper bounds on
the Lorentz-violating (LV) coefficients \cite{RMP.2011.83.11}.
The physical properties of the physical systems can be accessed by
including in all sectors of the minimal standard model LV terms.
In the fermion sector, for example, this violation is
implemented by introducing two CPT-odd terms,
$V_{\mu}\bar{\psi}\gamma ^{\mu }\psi $, $W_{\mu}\bar{\psi}
\gamma_{5}\gamma ^{\mu }\psi $, where $V_{\mu}$, $W_{\mu}$
are the LV backgrounds.
The LV terms are generated as vacuum expectation values of
tensors defined in a high energy scale.
The SME has also been used as a
framework to propose Lorentz violation
\cite{PRL.2001.87.251304,PRD.2002.66.056005} and CPT
\cite{PRL.1997.79.1432,PRD.1998.57.3932,PRL.2002.88.090801,
PRL.2000.84.1381,PRL.2000.84.1098,PRL.1999.82.2254} probing
experiments, which have amounted to the imposition of stringent
bounds on the LV coefficients.
By carefully analyzing the sectors of the SME some authors have
specialized in introducing news nonminimal couplings between
fermionic and gauge fields in the context of the Dirac equation.
In Ref. \cite{EPJC.2005.41.421}, for example, a LV
and CPT-odd nonminimal coupling between fermions and the gauge
field was proposed in the form
\begin{equation}
  D_{\mu}=\partial_{\mu}+ieA_{\mu}+i\frac{g}{2}
  \epsilon_{\mu\lambda\alpha\beta}\upsilon^{\lambda}F^{\alpha\beta},
  \label{eq:cov}
\end{equation}
in the context of the Dirac equation,
\begin{equation}
  (i\gamma^{\mu}D_{\mu}-M)\Psi=0,
  \label{eq:dirac1}
\end{equation}
were $\Psi $ is the fermion spinor,
$\upsilon^{\mu}=(\upsilon_{0},\boldsymbol{\upsilon})$ is the
Carroll-Field-Jackiw four-vector, $g$ is a constant that
measures the nonminimal coupling magnitude, and $F^{\mu\nu}$
is the electromagnetic field tensor, with
\begin{equation}
  F^{0i}=-F^{i0}=E^{i},~~~F^{ij}=-F^{ji}=\epsilon ^{ijk}\,B_{k}.
  \label{eq:fields}
\end{equation}
This suggests that the LV background, intervening in spacetime,
may correct or generate some new properties to the particles.
The analysis of the nonrelativistic limit of
Eq. \eqref{eq:dirac1} reveals that the nonminimal coupling of the
background with the electromagnetic fields generates a magnetic
dipole moment $g\boldsymbol{\upsilon}$ even for non-charged particles
\cite{EPJC.2005.41.421}, yielding an Aharonov-Casher (AC) phase
\cite{PRL.1984.53.319} for its wave function.
The nonminimal coupling in Eq. \eqref{eq:cov} has been applied to
several physical systems in relativistic and nonrelativistic quantum
mechanics
\cite{PRD.2012.86.045001,ADP.2011.523.910,JMP.2011.52.063505,
PRD.2011.83.125025,PLB.2006.639.675,EPJC.2009.62.425,
JPG.2012.39.055004,JPG.2012.39.105004,AP.2013.333.272,
JPG.2012.39.085001,JPG.2013.40.065002}.
Recently, a new CPT-even and LV dimension-five nonminimal
coupling between fermionic and gauge fields, involving the
CPT-even and Lorentz-violating gauge tensor of the SME was
proposed in Ref. \cite{PRD.2013.87.047701}
(cf. also Ref. \cite{PRD.2012.86.125033}).
This new nonminimal coupling was identified by
\begin{equation}
  D_{\mu}=\partial_{\mu}+ieA_{\mu}+\frac{\lambda}{2}
  \left(K_{F}\right)_{\mu\nu\alpha\beta }
  \gamma^{\nu}F^{\alpha\beta},
  \label{eq:covader}
\end{equation}
with $\left(K_{F}\right)_{\mu\nu\alpha\beta}$ being the
tensor ruling  the Lorentz violation in the CPT-even
electrodynamics of the SME.
This nonminimal coupling modifies the Dirac equation, whose
nonrelativistic regime is governed by a Hamiltonian which
induces new effects, such as an electric-Zeeman-like spectrum
splitting and an anomalous-like contribution to the electron
magnetic moment, among others.

In this paper, we specialize to the nonminimal coupling in
Eq. \eqref{eq:cov}.
The aim is to study the effects of this LV background in the scattering
process of a spin-1/2 neutral particle with magnetic dipole moment
$g\boldsymbol{\upsilon}$ in the presence of a electric field of an
infinitely long, infinitesimally thin line of charge, 
in the nonrelativistic limit.

The work is outlined in the following way: In Section
\ref{sec:motion} we derive the Schr\"{o}dinger-Pauli equation in
order to study the physical implications of the LV background on
the spin-1/2 AC scattering problem.
The Section \ref{sec:selfae} is devoted to the study of the LV
Hamiltonian via the self-adjoint extension technique and are
presented some important properties of the LV wave function.
In Section \ref{sec:scattbound} are addressed the scattering and
bound-state problems within the framework of the LV
Schr\"{o}dinger-Pauli equation.
Expressions for the energy bound-states, phase-shift and the
scattering matrix are computed and all them are explicitly
described in terms of the physical condition of the problem and,
as it was expected, the self-adjoint extension parameter is also
expressed in terms of the physical parameters.
At the end, we make a detailed analysis of the helicity
conservation's problem in the present framework.
In Section \ref{sec:conclusion} we give our conclusions and
remarks.

\section{The equation of motion}
\label{sec:motion}

In this section, we derive the equation of motion that governs
the dynamics of a spin-1/2 neutral particle in a radial electric
field and a LV background.
We start with the (2+1)-dimensional Dirac equation, which
follows from the decoupling of (3+1)-dimensional Dirac equation
for the specialized case where $\partial _{3}=0$, into two
uncoupled two-component equations, such as implemented in
Refs. \cite{PRD.1978.18.2932,NPB.1988.307.909,PRL.1989.62.1071}.
Since we are interested only in the effects of the LV
background, we can consider only the sector generating the AC
effect in Eq. \eqref{eq:cov}.
In this case, the planar Dirac equation ($\hbar=c=1$) is
\begin{equation}
  \left(
    \beta \mathbf{\gamma }\cdot \boldsymbol{\Pi}+\beta M
  \right) \Psi =
  \bar{E} \Psi,
\label{eq:defdirac}
\end{equation}
where $\Psi$ is a two-component spinor,
$\boldsymbol{\Pi}=\mathbf{p}-
gs\left(\boldsymbol{\upsilon}\times \mathbf{E}\right)$
is the generalized momentum, and $s$ is twice the spin value,
with $s=+1$ for spin ``up'' and $s=-1$ for spin ``down''.
The $\gamma$-matrices in $(2+1)$ dimensions are given in terms of the
Pauli matrices
\begin{equation}
  \beta = \gamma _{0}=\sigma _{3},\qquad
  \gamma_{1}=i\sigma _{2},\qquad
\gamma_{2}=-is\sigma _{1}.
\end{equation}
The field configuration (in cylindrical coordinates) is chosen
to be
\begin{equation}
  \mathbf{E}=2\lambda
  \frac{\mathbf{\hat{r}}}{r},\qquad
  \boldsymbol{\nabla }\cdot\mathbf{E}=
  2\lambda\frac{\delta(r)}{r},\qquad
  \upsilon^{\mu}=(0,0,0,\upsilon),
  \label{eq:fieldac}
\end{equation}
where $\mathbf{E}$ is the electric field generated by an infinite
charge filament and $\lambda$ is the linear charge density along the
$z$-axis.
The second order equation implied by Eq. \eqref{eq:defdirac} is
obtained by applying the matrix operator
$\left[M+\beta\bar{E}-
\mathbf{\gamma}\cdot\boldsymbol{\Pi}\right]\beta$.
After this application, one finds
\begin{align}
  \left(\bar{E}^{2}-M^{2}\right)\Psi
  = {} & -
  \left(\mathbf{\gamma}\cdot\boldsymbol{\Pi}\right)
  \left(\mathbf{\gamma}\cdot\boldsymbol{\Pi}\right)
  \Psi \nonumber \\
  = {} &
  \left\{
    \boldsymbol{\Pi}^{2}+g\mathbf{\sigma }\cdot
    \left[
      \boldsymbol{\nabla}\times
      \left(\boldsymbol{\upsilon}\times \mathbf{E}\right)
    \right]
  \right\} \Psi.
\label{eq:hrel}
\end{align}
By accessing the nonrelativistic limit,
$\bar{E}=M+E$, $M\gg E$, we obtain the
Schr\"{o}dinger-Pauli equation
\begin{equation}
  \hat{H}\Psi =E\Psi,
  \label{eq:dedfm}
\end{equation}
with
\begin{equation}
  \hat{H}=\frac{1}{2M}
  \left[
    \mathbf{p}-gs\left(\boldsymbol{\upsilon}\times \mathbf{E}\right)
  \right]^{2}+
  \frac{1}{2M}g\mathbf{\sigma}\cdot
  \left[
    \boldsymbol{\nabla }\times(\boldsymbol{\upsilon}\times \mathbf{E})
  \right],
  \label{eq:hdef}
\end{equation}
the Hamiltonian operator.
Using \eqref{eq:fieldac}, the Hamiltonian \eqref{eq:hdef}
becomes
\begin{equation}
  \hat{H}=
  \frac{1}{2M}
  \left[
  \hat{H}_{0} +
  \alpha \sigma_{z} \frac{\delta(r)}{r}
  \right],
  \label{eq:hnrf}
\end{equation}
with
\begin{equation}
    \label{eq:hzero}
  \hat{H}_{0}=
  \left(
    \frac{1}{i}\boldsymbol{\nabla}
    -\alpha s\frac{\hat{\boldsymbol{\varphi}}}{r}
  \right)^{2},
\end{equation}
and
\begin{equation}
\alpha = 2g \upsilon\lambda,
\label{eq:deltac}
\end{equation}
is the coupling constant of the $\delta(r)/r$ potential.

The Hamiltonian in Eq. \eqref{eq:hnrf} governs the quantum dynamics of a
spin-1/2 neutral particle with a radial electric field, i.e., a spin-1/2
AC problem, with $g\boldsymbol{\upsilon}$ playing the role of a
nontrivial magnetic dipole moment, in contrast with the usual AC problem
where the magnetic dipole moment is
$\boldsymbol{\mu}=\mu\sigma_{z} \hat{\mathbf{z}}$
\cite{EPJC.2013.73.2402}.
Also, in Eq. \eqref{eq:hnrf} we observe the presence of a $\delta$
function which is singular at the origin.
This makes the problem more complicated to be solved.
Such kind of point interaction potential can then be addressed
by the self-adjoint extension approach
\cite{JMP.1985.26.2520,Book.2004.Albeverio}, which will be used
for studying the scattering and bound state scenarios.

\section{Self-adjoint extension analysis}
\label{sec:selfae}

An operator $\mathcal{O}$, with domain
$\mathcal{D}(\mathcal{O})$, is said to be self-adjoint if and
only if
$\mathcal{D}(\mathcal{O}^{\dagger})=\mathcal{D}(\mathcal{O})$
and $\mathcal{O}^{\dagger}=\mathcal{O}$.
In order to determine all self-adjoint extensions of
\eqref{eq:hzero}, making use of the underlying rotational
symmetry expressed by the fact that $[\hat{H},\hat{J}_{z}]=0$,
where $\hat{J}_{z}=-i\partial/\partial_{\varphi}+\sigma_{z}/2$ is
the total angular momentum operator in the $z$-direction,
we decompose the Hilbert space
$\mathfrak{H}=L^{2}(\mathbb{R}^{2})$
with respect to the total angular momentum
$\mathfrak{H}=\mathfrak{H}_{r}\otimes\mathfrak{H}_{\varphi}$,
where $\mathfrak{H}_{r}=L^{2}(\mathbb{R}^{+},rdr)$ and
$\mathfrak{H}_{\varphi}=L^{2}(\mathcal{S}^{1},d\varphi)$,
with $\mathcal{S}^{1}$ denoting the unit sphere in
$\mathbb{R}^{2}$.
So, it is possible to express the eigenfunctions of the two
dimensional Hamiltonian in terms of the eigenfunctions of
$\hat{J}_{z}$
\begin{equation}
  \Psi(r,\varphi)=
  \left(
    \begin{array}{c}
      \psi_{m}(r) e^{i(m_{j}-1/2)\varphi } \\
      \chi_{m}(r) e^{i(m_{j}+1/2)\varphi }
    \end{array}
  \right) ,
\label{eq:wavef}
\end{equation}
with $m_{j}=m+1/2=\pm 1/2,\pm 3/2,\ldots $, and $m\in
\mathbb{Z}$.
By inserting Eq. \eqref{eq:wavef} into Eq. \eqref{eq:dedfm} the
Schr\"{o}dinger-Pauli equation for $\psi_{m}(r)$ is found to be
($k^{2}=2ME$)
\begin{equation}
  H\psi_{m}(r)=k^{2}\psi_{m}(r),
  \label{eq:eigen}
\end{equation}
where
\begin{equation}
  H=H_{0}+\alpha \frac{\delta(r)}{r},
  \label{eq:hfull}
\end{equation}
and
\begin{equation}
  H_{0}=
  -\frac{d^{2}}{dr^{2}}-\frac{1}{r}\frac{d}{dr}
  +\frac{(m-\alpha s)^{2}}{r^{2}}.
\end{equation}

The self-adjoint extension approach consists, essentially, in
extending the domain $\mathcal{D}(H_{0})$ to match
$\mathcal{D}(H_{0}^{\dagger})$ and therefore turning $H_{0}$
into a self-adjoint operator.
To do so, we must find the deficiency subspaces,
$N_{\pm}$, with dimensions $n_{\pm}$, which are called
deficiency indices of $H_{0}$ \cite{Book.1975.Reed.II}.
A necessary and sufficient condition for $H_{0}$ being
essentially self-adjoint is that $n_{+}=n_{-}=0$.
On the other hand, if $n_{+}=n_{-}\geq 1$, then $H_{0}$ has an
infinite number of self-adjoint extensions parametrized by
unitary operators  $U:N_{+}\to N_{-}$.
In order to find the deficiency subspaces of $H_{0}$ in
$\mathfrak{H}_{r}$, we must solve the eigenvalue equation
\begin{equation}
  H_{0}^{\dagger}\psi_{\pm}
  =\pm i k_{0}^{2} \psi_{\pm},
  \label{eq:eigendefs}
\end{equation}
where $k_{0}^{2}\in\mathbb{R}$ was introduced for dimensional
reasons.
Since $H_{0}^{\dagger}=H_{0}$, the solutions of
Eq. \eqref{eq:eigendefs} which vanishes at the infinite are the
Hankel functions (up to a constant)
\begin{equation}
\psi_{\pm}=H_{|m-\alpha s|}^{(1)}(\sqrt{\mp i} k_{0} r),
\label{eq:psidefs}
\end{equation}
with $\rm{Im} \sqrt{\pm i}>0$.
The dimension of such deficiency subspace is thus
$(n_{+},n_{-})=(1,1)$.
According to the von Neumann-Krein theory, all self-adjoint
extensions $H_{\theta,0}$ of $H_{0}$ are given by the
one-parameter family
\begin{equation}
\mathcal{D}(H_{\theta,0})=\mathcal{D}(H_{0}^{\dagger})=
\mathcal{D}(H_{0})\oplus N_{+}\oplus N_{-}.
\end{equation}
Thus, $\mathcal{D}(H_{\theta,0})$ in $\mathfrak{H}_{r}$ is given
by the set of functions \cite{Book.1975.Reed.II}
\begin{equation}
  \label{eq:domain}
  \psi_{\theta}(r) = \psi_{m}(r) +
  c\left[ H_{|m-\alpha s|}^{(1)}(\sqrt{-i} k_{0}r)+
    e^{i\theta}H_{|m-\alpha s|}^{(1)}(\sqrt{i}k_{0}r) \right] ,
\end{equation}
where $\psi_{m}(r)$, with $\psi_{m}(0)=\dot{\psi}_{m}(0)=0$
($\dot{\psi}\equiv d\psi/dr$), is the regular wave function,
$c\in \mathbb{C}$ and the number $\theta \in [0,2\pi)$
represents a choice for the boundary condition.
Using the unitary operator
$U: L^{2}(\mathbb{R}^{+},rdr) \to L^{2}(\mathbb{R}^{+}, dr)$,
given by $(U \xi)(r)=r^{1/2}\xi(r)$, the operator $H_{0}$
becomes
\begin{equation}
  \tilde{H}_{0}=
  U H_{0} U^{-1}=
  -\frac{d^{2}}{dr^{2}}
  +\frac{(m-\alpha s)^{2}-1/4}{r^{2}}.
\end{equation}
By standard results the radial operator $\tilde{H}_{0}$,
is essentially self-adjoint for $|m-\alpha s| \geq 1$, while
for $|m-\alpha s|< 1$ it admits an one-parameter family of
self-adjoint extensions \cite{Book.1975.Reed.II}.
This statement can be understood based in
Eq. \eqref{eq:psidefs}, because for $|m-\alpha s| \geq 1$ the
right hand side is not in $\mathfrak{H}_{r}$ at $0$, while it is
in  $\mathfrak{H}_{r}$ for $|m-\alpha s|<1$.

All the self-adjoint extensions $H_{0,\lambda_{m}}$ of
$\tilde{H}_{0}$ are parametrized by the boundary condition at
the origin \cite{JMP.1985.26.2520,Book.2004.Albeverio}
\begin{equation}
\psi^{(0)}=\lambda_{m}\psi^{(1)},  \label{eq:bc}
\end{equation}
with
\begin{align*}
  \psi^{(0)}={} & \lim_{r\rightarrow 0^{+}}r^{|m-\alpha s|}\psi_{m}(r), \\
  \psi^{(1)}={} & \lim_{r\rightarrow 0^{+}}\frac{1}{r^{|m-\alpha s|}}
  \left[\psi_{m}(r)-\psi^{(0)}\frac{1}{r^{|m-\alpha s|}}\right],
\end{align*}
where $\lambda_{m}$ is the self-adjoint extension parameter.
In \cite {Book.2004.Albeverio} is shown that there is a
relation between the self-adjoint extension parameter
$\lambda_{m}$ and the number $\theta$ in Eq. \eqref{eq:domain}.
The number $\theta$ is associated with the mapping of
deficiency subspaces and extend the domain of operator to make
it self-adjoint.
The self-adjoint extension parameter $\lambda_{m}$ have a
physical interpretation: it represents the scattering length
\cite{Book.2011.Sakurai} of $H_{0,\lambda_{m}}$
\cite{Book.2004.Albeverio}.
For $\lambda_{m}=0$ we have the free Hamiltonian (without the
$\delta$ function) with regular wave functions at origin and
for $\lambda_{m}\neq 0$ the boundary condition in Eq. \eqref{eq:bc}
allows a $r^{-|m-\alpha s|}$ singularity in the wave functions at
origin.

\section{Scattering and bound state analysis}
\label{sec:scattbound}

The general solution for Eq. \eqref{eq:eigen} in the $r\neq 0$
region can be written as
\begin{equation}
\label{eq:sol1}
\psi_{m}(r)=a_{m}J_{|m-\alpha s|}(kr)+b_{m}Y_{|m-\alpha s|}(kr),
\end{equation}
with $a_{m}$ and $b_{m}$ being constants and $J_{\nu}(z)$ and
$Y_{\nu}(z)$ are the Bessel functions of first and second kind,
respectively.
Upon replacing $\psi_{m}(r)$ in the boundary condition
\eqref{eq:bc}, one obtain
\begin{equation}
  \lambda_{m} a_{m} \mathcal{A} k^{|m-\alpha s|}=
  b_{m}\bigg[\mathcal{B} k^{-|m-\alpha s|}
  -\lambda_{m}\Big(\mathcal{C}  k^{|m-\alpha s|}
  + \mathcal{B} \mathcal{D} k^{-|m-\alpha s|}
  \lim_{r\rightarrow 0^{+}}r^{2-2|m-\alpha s|}\Big)\bigg],
  \label{eq:bcf}
\end{equation}
with
\begin{align}
\mathcal{A} & = \frac{1}{2^{|m-\alpha s|}\Gamma(1+|m-\alpha
  s|)}, &
\mathcal{B} & =-\frac{2^{|m-\alpha s|}\Gamma(|m-\alpha s|)}{\pi},
\nonumber \\
\mathcal{C} &=-\frac{\cos (\pi |m-\alpha s|)
 \Gamma(-|m-\alpha s|)}{\pi 2^{|m-\alpha s|}}, &
\mathcal{D} &= \frac{k^{2}}{4(1-|m-\alpha s|)}.
\end{align}
In Eq. \eqref{eq:bcf}, $\lim_{r\rightarrow 0^{+}}r^{2-2|m-\alpha s|}$
is divergent if $|m-\alpha s|\geq 1$, hence $b_{m}$ must be zero.
On the other hand, $\lim_{r\rightarrow 0^{+}}r^{2-2|m-\alpha s|}$
is finite for $|m-\alpha s|<1$.
This means that there arises the contribution of the irregular
solution $Y_{|m-\alpha s|}(kr)$.
Here, the presence of an irregular solution contributing to the
wave function stems from the fact the Hamiltonian $H_{0}$ is not a
self-adjoint operator when $|m-\alpha s|<1$ (cf. Section
\ref{sec:selfae}), hence such irregular solution must be
associated with a self-adjoint extension of the operator $H_{0}$
\cite{JPA.1995.28.2359,PRA.1992.46.6052}.
Thus, for $|m-\alpha s|<1$, we have
\begin{equation}
  \lambda_{m} a_{m}\mathcal{A} k^{|m-\alpha s|}=
  b_{m}(\mathcal{B} k^{-|m-\alpha s|}-
  \lambda_{m}\mathcal{C} k^{|m-\alpha s|}),
\end{equation}
and by substituting the values of $\mathcal{A}$, $\mathcal{B}$
and $\mathcal{C}$ into above expression we find
\begin{equation}
  b_{m}=-\mu_{m}^{\lambda_{m}}(k,\alpha) a_{m},
\end{equation}
where
\begin{equation}
  \mu_{m}^{\lambda_{m}}(k,\alpha)=
  \frac
  {\lambda_{m} k^{2|m-\alpha s|}
    \Gamma(1-|m-\alpha s|)\sin(\pi |m-\alpha s|)}
  {\lambda_{m} k^{2|m-\alpha s|}
  \Gamma{(1-|m-\alpha s|)}\cos(\pi|m-\alpha s|)+
  4^{|m-\alpha s|}\Gamma(1+|m-\alpha s|)}.
  \label{eq:mul}
\end{equation}
Since a $\delta$ function is a very short range potential, it
follows that the asymptotic behavior of $\psi_{m}(r)$ for
$r\rightarrow \infty$ is given by \cite{JPA.2010.43.354011}
\begin{equation}
  \psi_{m}(r)\sim \sqrt{\frac{2}{\pi kr}}
  \cos \left[ kr-\frac{|m|\pi}{2}-
  \frac{\pi}{4}+
  \delta_{m}^{{\lambda_{m}}}(k,\alpha)\right] ,
\label{eq:f1asim}
\end{equation}
where $\delta_{m}^{{\lambda_{m}}}(k,\alpha)$  is a
scattering phase shift.
The phase shift is a measure of the argument difference to the
asymptotic behavior of the solution $J_{|m|}(kr)$ of the radial
free equation which is regular at the origin.
By using the asymptotic behavior of the Bessel functions
\cite{Book.1972.Abramowitz} into Eq. \eqref{eq:sol1} one obtain
\begin{align}
  \label{eq:scattsol}
  \psi_{m}(r)
  \sim {}
  &
  a_{m}
  \sqrt{\frac{2}{\pi kr}}
  \left\{
    \cos\left[kr-\frac{\pi |m-\alpha s|}{2}-\frac{\pi}{4}\right]
  \right.
  \nonumber \\
  &
  \left.
    -\mu_{m}^{{\lambda_{m}}}(k,\alpha)
    \sin\left[ kr-\frac{\pi |m-\alpha s|}{2}-\frac{\pi}{4}\right]
  \right\} .
\end{align}
By comparing the above expression with Eq. \eqref{eq:f1asim}, we
have
\begin{equation}
  \cos
  \left[
    kr-\frac{\pi |m-\alpha s|}{2}-\frac{\pi}{4}+
    \theta_{m}^{\lambda_{m}}(k,\alpha)
  \right]=
  \cos
  \left[
    kr-\frac{\pi|m|}{2}-\frac{\pi}{4}+
    \delta_{m}^{{\lambda_{m}}}(k,\alpha)
  \right].
\end{equation}
with $\theta_{m}^{\lambda_{m}}(k,\alpha)$ given by
\begin{equation}
  \cos \left[\theta_{m}^{\lambda_{m}}(k,\alpha)\right] = a_{m}, \qquad
      \sin\left[\theta_{m}^{\lambda_{m}}(k,\alpha)\right] = a_{m}
    \mu_{m}^{{\lambda_{m}}}(k,\alpha).
    \label{eq:sincostheta}
\end{equation}
Thus, the asymptotic behavior in Eq. \eqref{eq:f1asim} is satisfied if
\begin{equation}
a_{m}=\frac{1}{\sqrt{1+\left[\mu_{m}^{{\lambda_{m}}}(k,\alpha)\right]^{2}}}.
\end{equation}
Now, comparing the arguments of the cosines above, the sought phase
shift is obtained
\begin{equation}
\delta_{m}^{{\lambda_{m}}}(k,\alpha)=
\Delta_{m}(\alpha)+\theta_{m}^{\lambda_{m}}(k,\alpha),
\label{eq:phaseshift}
\end{equation}
with
\begin{equation}
\Delta_{m}(\alpha)=\frac{\pi}{2}(|m|-|m-\alpha s|),
\end{equation}
the phase shift of the AC scattering and
\begin{equation}
  \theta_{m}^{\lambda_{m}}(k,\alpha)=\arctan {[\mu_{m}^{\lambda_{m}}(k,\alpha)]}.
\end{equation}
Therefore, by using Eq. \eqref{eq:sincostheta}, the scattering operator
$S_{m}^{\lambda_{m}}(k,\alpha)$ ($S$-matrix) for the self-adjoint
extension is
\begin{equation}
S_{m}^{\lambda_{m}}(k,\alpha)=
e^{2i\delta_{m}^{{\lambda_{m}}}(k,\alpha)}=
  \left[
    \frac
    {1+i\mu_{m}^{\lambda_{m}}(k,\alpha)}
    {1-i\mu_{m}^{\lambda_{m}}(k,\alpha)}
  \right]
  e^{2i\Delta_{m}(\alpha)},
\end{equation}
or by using Eq. \eqref{eq:mul}, we have
\begin{equation}
  S_{m}^{\lambda_{m}}(k,\alpha)
  =
  \left[
    \frac
    {\lambda_{m} k^{2|m-\alpha s|}
      \Gamma(1-|m-\alpha s|)e^{i\pi |m-\alpha s|}+
      4^{|m-\alpha s|}\Gamma(1+|m-\alpha s|)
    }
    {\lambda_{m} k^{2|m-\alpha s|}
      \Gamma(1-|m-\alpha s|)e^{-i\pi |m-\alpha s|}+
      4^{|m-\alpha s|}\Gamma(1+|m-\alpha s|)
    }
  \right]
  e^{2i\Delta_{m}(\alpha)}.
  \label{eq:smatrix}
\end{equation}
Hence, for any value of the self-adjoint extension parameter
$\lambda_{m}$, there is an additional scattering.
If ${\lambda_{m}}=0$, we achieve the corresponding result
for the AC problem with Dirichlet boundary condition
\cite{AP.1983.146.1}, $S_{m}^{0}(k,\alpha)=e^{2i\Delta_{m}(\alpha)}$.
For ${\lambda_{m}}=\infty$, we get
$S_{m}^{\infty}(k,\alpha)=e^{2i\Delta_{m}(\alpha)+2i\pi |m-\alpha s|}$.

In accordance with the general theory of scattering, the poles
of the $S$-matrix in the upper half of the complex plane
\cite{PRC.1999.60.34308} determine the positions of the bound
states in the energy scale.
These poles occur in the denominator of Eq. \eqref{eq:smatrix} with
the replacement $k\rightarrow i\kappa$, with $\kappa^{2}=-2ME$,
$E<0$.
Thus,
\begin{equation}
  \lambda_{m} (i\kappa)^{2|m-\alpha s|}
  \Gamma(1-|m-\alpha s|)e^{-i\pi |m-\alpha s|}+
  4^{|m-\alpha s|}\Gamma(1+|m-\alpha s|)=0.
\end{equation}
Solving the above equation for $E$, we found the bound state
energy
\begin{equation}
  E=-\frac{2}{M}
  \left[-\frac{1}{\lambda_{m}}
    \frac{\Gamma(1+|m-\alpha s|)}{\Gamma(1-|m-\alpha s|)}
  \right]^{1/|m-\alpha s|},
  \label{eq:energy_BG-sc}
\end{equation}
for $\lambda_{m}<0$.
Hence, the poles of the scattering matrix only occur for
negative values of the self-adjoint extension parameter, when we
have scattering and bound states.
In this latter case, the scattering operator can be expressed in
terms of the bound state energy
\begin{equation}
  S_{m}^{\lambda_{m}}(k,\alpha)=e^{2i\Delta_{m}(\alpha)}
  \left[
    \frac
    {e^{2 i \pi |m-\alpha s|}-(\kappa/k)^{2|m-\alpha s|}}
    {1-(\kappa/k)^{2|m-\alpha s|}}
  \right].
\end{equation}

In a previous work \cite{EPL.2013.101.51005} (cf. also
\cite{AP.2010.325.2529,PLB.2013.719.467,JMP.2012.53.122106}
for analogous systems), using another self-adjoint
extension approach, the energy bound state for the present
system was determined in terms of the physics of the problem,
and it read
\begin{equation}
  E=
  -\frac{2}{M a^{2}}
  \left[
    \left(
      \frac
      {\alpha +|m-\alpha s|}
      {\alpha -|m-\alpha s|}
    \right)
    \frac
    {\Gamma (1+|m-\alpha s|)}
    {\Gamma (1-|m-\alpha s|)}
  \right]^{{1}/{|m-\alpha s|}},
\label{eq:energy_KS}
\end{equation}
where $a$ is a very small radius smaller than the Compton
wavelength $\lambda_{C}$ of the electron
\cite{PLB.1994.333.238}, which comes from the
regularization of the $\delta$ function.
By comparing Eq. \eqref{eq:energy_BG-sc} with Eq. \eqref{eq:energy_KS}
we have
\begin{equation}
  \frac{1}{\lambda_{m}}=
  -\frac{1}{a^{2|m-\alpha s|}}
    \left(
      \frac
      {\alpha +|m-\alpha s|}
      {\alpha -|m-\alpha s|}
    \right).
\label{eq:saep}
\end{equation}
The above relation is only valid for $\lambda_{m}<0$,
consequently we have $|\alpha|\geq |m-\alpha s|$ and due to
$|m-\alpha s|<1$ it is sufficient to consider $|\alpha|\geq 1$ to
guarantee $\lambda_{m}$ to be negative.
A necessary condition for a $\delta$ function generates an
attractive potential, which is able to support bound states, is
that the coupling constant must be negative.
Thus, the existence of bound states requires
\begin{equation}
\alpha \leq -1.
\end{equation}
Also, it seems from the above equation and from Eq. \eqref{eq:deltac}
that we must have $g\upsilon\lambda < 0$ and there is a minimum value
for this product to ensure the presence of a bound state.

The scattering amplitude $f(k,\alpha)$ can be now obtained using
the standard methods of scattering theory, namely
\begin{align}
  f(k,\alpha)
  = {} &
  \frac{1}{\sqrt{2\pi i k}}\sum_{m=-\infty}^{\infty}
  \left(
    S_{m}^{\lambda_{m}}(k,\alpha)-1
  \right)
  e^{im\varphi}   \nonumber \\
  ={} &
    \frac{1}{\sqrt{2\pi i k}}
    \left\{
      \sum_{|m-\alpha s|\geq 1}
      (e^{2 i \Delta_{m}(\alpha)}-1)e^{im\varphi}
    \right. \nonumber \\
    & +
    \left.
    \sum_{|m-\alpha s|< 1}
    (e^{2 i \Delta_{m}(\alpha)}
    \left[
      \frac
      {1+ i\mu_{m}^{\lambda_{m}}(k,\alpha)}
      {1-i\mu_{m}^{\lambda_{m}}(k,\alpha)}
    \right]-1)e^{im\varphi}
  \right\}.
  \label{eq:scattamp}
\end{align}
The first sum is the AC amplitude
(i.e., when the $\delta$ function is absent),
while the second sum is the contribution that come from the singular
solutions.
In the above equation we can see that the scattering amplitude
is energy dependent (cf. Eq. \eqref{eq:mul}).
This is a clearly manifestation of the known non-conservation of
the helicity in the AC scattering \cite{PRL.1990.64.2347},
because the only length scale in the nonrelativistic problem is
set by  $1/k$, so it follows that the scattering amplitude would
be a function of the angle alone, multiplied by $1/k$
\cite{PRD.1977.16.1815}.
In fact, the failure of helicity conservation expressed in
Eq. \eqref{eq:scattamp}, it stems from the fact that the
$\delta$ function singularity make the Hamiltonian and the
helicity nonself-adjoint operators
\cite{PRD.1977.15.2287,PRL.1983.50.464,PLB.1993.298.63,
NPB.1994.419.323}.
By expressing the helicity operator,
$\hat{h}=\boldsymbol{\Sigma} \cdot \boldsymbol{\Pi}$,
in terms of the variables used in Eq. \eqref{eq:wavef}, we attain
\begin{equation}
  \hat{h} =
  \left(
    \begin{array}{cc}
      0
      & \displaystyle -i
      \left(
        \partial_{r}+\frac{|m-\alpha s|+1}{r}
      \right) \\
      \displaystyle -i
      \left(
        \partial_{r}-\frac{|m-\alpha s|}{r}
      \right)
      & 0
    \end{array}
  \right).
\end{equation}
Notice under a parity $\pi$ transformation
$\hat{h}\to\pi^{\dagger}\hat{h}\pi=-\hat{h}$, that comes
immediately from the parity transformation
$\pi^{\dagger}r\pi=-r$.
This is in fact the helicity odd-parity property.
The helicity operator share the same issue as the Hamiltonian
operator in the interval $|m+\alpha s|<1$, i.e., it is not
self-adjoint \cite{PRD.1994.49.2092,JPA.2001.34.8859}.
Despite that on a finite interval $[0,L]$, $\hat{h}$  is a
self-adjoint operator with domain in the functions satisfying
$\xi(L)=e^{i\theta}\xi{(0)}$, it does not admit a self-adjoint
extension on the interval $[0,\infty)$ \cite{AJP.2001.69.322},
and consequently it cannot be conserved, thus the helicity
conservation is broken due to the presence of the singularity at
the origin \cite{PRD.1977.16.1815,PRL.1983.50.464}.

\section{Conclusion}
\label{sec:conclusion}

We have studied the spin-1/2 AC scattering problem with a
Lorentz-violating and CPT-odd nonminimal coupling between
fermions and the gauge field in the context of the 
nonrelativistic limit of the Dirac equation.
It has been shown that there is an additional scattering for any
value of the self-adjoint extension parameter and for negative
values of this parameter there are non-zero energy bound states.
The scattering amplitude show a energy dependency, so the
helicity in not conserved.
This stem from the fact that the helicity operator is not
a self-adjoint extension operator.

\section*{Acknowledgments}
F. M. Andrade acknowleges reasearcher grants by Funda\c{c}\~{a}o
Arauc\'{a}ria project No. 205/2013 and 
E. O. Silva acknowledges researcher grants by CNPq-(Universal) project
No. 484959/2011-5.

\section*{References}
\bibliographystyle{unsrt}

\end{document}